OpenStudy (anonymous):

]Which of the following is the slope between the points (0, -10) and (-4, 0)?

OpenStudy (lgbasallote):

\[\huge slope = \frac{y_2 - y_1}{x_2 - x_1}\] \[\huge slope = \frac{0 - (-10)}{-4 - 0}\]

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

so what would be the answer? lol

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

2/5 ?

OpenStudy (anonymous):

Slope is \[DeltaY/DeltaX\] so slope is \[0-(-10)\div(-4)-0=5\div2\]

OpenStudy (anonymous):

so it is -5/2 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you! :)